OpenStudy (anonymous):

If (x – 5) is a factor of f(x), which of the following must be true? A root of f(x) is x = –5. B) A root of f(x) is x = 5. C) Both x = –5 and x = 5 are roots of f(x). D) Neither x = –5 nor x = 5 is a root of f(x).

OpenStudy (anonymous):

if\[ f(x)=(x-5)\times \text{something}\] then \[f(5)=(5-5)\times \text{something}\] i.e. \[f(5)=0\]

OpenStudy (anonymous):

since \(f(5)=0\) we say " a root of \(f(x)\) is \(5\)"

OpenStudy (anonymous):

it'd B

OpenStudy (anonymous):

@satellite73